# Information Elicitation in Agency Games

### v2

@@ A.3 Comparison of decreasing ratio condition to increasing ratio condition
A.4 Proofs from Section 3.4
Lemmas 2 and 3 show that the principal always benefits from more information being revealed. Assumption 3 further implies that the principal strictly benefits from revelation.
@@
A.5 Proofs from Section 3.5
@@
Total welfare increases under revelation ($W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})$) only if task completion quantity increases under revelation,
$F(p^{*})<(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*}).$
Proof.
[…]
Upper bounding this:
[…]
The right hand term is exactly the difference in output:
[…]
$=(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-((1-\theta)F_{0}(p^{*})+\theta F_{1}(p^{*}))$
$=(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-F(p^{*})$
$\implies W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})\leq(b-p^{*})\left((1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-F(p^{*})\right)$
Since $(b-p^{*})>0$, if $(1-\theta)F_{0}(p_{0}^{*})+\theta F_{1}(p_{1}^{*})-F(p^{*})\leq 0$, then $W_{\text{rev}}(p_{0}^{*},p_{1}^{*})-W_{\text{con}}(p^{*})\leq 0$. ∎
Appendix B Proofs from Section 4
B.1 Proofs from Section 4.2
@@ Appendix B Proofs from Section 4
Suppose $\gamma=\theta=\frac{1}{2}$. Suppose $F_{0},F_{1}$ are continuously differentiable CDFs, the principal’s utility is strictly concave (Assumption 1), and the MLRP holds (Assumption 3). Then $p_{0}^{\prime}(\varepsilon)>0$ and $p_{1}^{\prime}(\varepsilon)<0$ for all $\varepsilon\in[0,1]$.
@@
B.2 Proofs from Section 4.3
We first expand $V_{\text{garb}}(\varepsilon)$ for $\gamma=\theta=\frac{1}{2}$.
@@
$+\frac{1}{2}\left((V_{0}(p_{0}(\varepsilon))-V_{0}(p_{1}(\varepsilon)))-(V_{1}(p_{0}(\varepsilon))-V_{1}(p_{1}(\varepsilon)))\right)$
@@
$+\frac{1}{2}\left(-\Delta(p_{0}(\varepsilon),p_{1}(\varepsilon))\right)$
@@
$-\Delta(p_{0}(\varepsilon),p_{1}(\varepsilon))$
@@
$z_{x}(p)=\frac{V^{\prime}_{x}(p)}{\Pi_{x}^{{}^{\prime\prime}}(p)}$
@@
B.3 Proofs from Section 4.5
@@
$+\frac{1}{2}(b-p_{1}(\varepsilon))\left(\frac{1+\varepsilon}{2}F_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))\right)$
@@
By the MLRP and strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$, we have that $p_{1}(\varepsilon)<p_{0}(\varepsilon)<p_{0}^{*}$.
Strict concavity of $\Pi_{0}(p)$ and the optimality of $p_{0}^{*}$ for $\Pi_{1}$ then implies that
@@
B.4 Proofs from Section 4.6
Lemma 8 (Reducing garbling initially increases total welfare).
Proof.
Lemma 9 (Optimal garbling increases welfare over concealment).
Proof.
Proof.

### v3

@@ F.3 Comparison of Decreasing Ratio Condition to Increasing Ratio Condition
F.4 Proofs from Section D.4
Lemmas 8 and 9 show that the principal always benefits from more information being revealed. Assumption 6 further implies that the principal strictly benefits from revelation.
@@
Appendix G Additional Results on Information Revelation with Garbling
G.1 Garbling and Prices
Lemma 11 (Monotonic price changes).
Suppose $\theta=\frac{1}{2}$. Suppose the principal’s utility is strictly concave as a function of price (Assumption 4). Suppose the MLRP holds (Assumption 6). Then $p_{0}^{\prime}(\varepsilon)>0$ and $p_{1}^{\prime}(\varepsilon)<0$ for all $\varepsilon\in[0,1]$.
G.2 Agent’s Garbling Incentives: General Conditions
G.2.1 Garbling condition under one zero-cost type
As in Section D.2.1, we first analyze the restricted case where one of the agent types is anchored at zero-cost: suppose $C|X=1$ takes value 00 with probability one. Proposition 4 gives a sufficient condition for the agent to prefer a non-zero amount of garbling over full revelation.
Let the conditional distribution of the cost $C$ given $Y$ be distributed with CDF ${\mathbb{P}}(C\leq c|Y=y)=G_{y}(c)$. When $\theta=\frac{1}{2}$,
$G_{0}(c)=\frac{1+\varepsilon}{2}F_{0}(c)+\frac{1-\varepsilon}{2}F_{1}(c);\;\;G_{1}(c)=\frac{1+\varepsilon}{2}F_{1}(c)+\frac{1-\varepsilon}{2}F_{0}(c).$
Let $g_{y}(c)$ be denote the PDF.
Proposition 4 (Sufficient garbling condition with zero-cost type).
$\frac{b-p_{0}^{*}}{2-\sigma_{0}(p_{0}^{*})}<g_{0}(p_{0}^{*}),$ (19)
[…]
Notably, the inequality in equation (19) captures distributions that are not captured by Proposition 1. Thus, comparing Proposition 4 to Proposition 1 shows that the agent might want to garble, even if they may not always want to hide. In fact, the condition in equation (19) is quite general, and we show in the example in Section G.3 below that equation (19) applies to any log-concave Weibull distribution.
G.2.2 General garbling condition
Definition 2 (Agent utility dominance).
Proposition 5 (Sufficient garbling condition).
G.3 Example: Exponential and Weibull Distributions
The significance of this example is that if one agent type is anchored at 00, and the non-zero-cost environment induces an exponential distribution, the agent will always have an incentive to garble, regardless of the mean of the non-zero-cost distribution. Consider this in comparison to the exponential example from Section D.3, where the condition in Proposition 1 showed that agent prefers to fully conceal $X$ when $\lambda_{0}$ is small enough.
Similarly to Figure 6, we can also consider simulations beyond the zero-cost anchored setting by considering all combinations of $\lambda_{0},\lambda_{1}$ when $C|X=0\sim\text{Exp}(\frac{1}{\lambda_{0}})$ and $C|X=1\sim\text{Exp}(\frac{1}{\lambda_{1}})$, with $F_{x}$ given by equation (12). Figure 7 illustrates the combinations of $\lambda_{0},\lambda_{1}$ for which the agent prefers some amount of garbling over full revelation. Figure 7 shows that $V_{\text{garb}}^{\prime}(1)<0$ as long one of the conditional means $\lambda_{0}$ or $\lambda_{1}$ is small enough. That is, as long as one of the revealed settings has low enough cost, the agent always prefers to garble, regardless of how high the cost of the other setting goes. This contrasts the revelation example in Section D.3, where even if $\lambda_{1}$ is close to 00, high enough $\lambda_{0}$ leads to the agent being willing to reveal.
{tblr}
p0.1cmc &$V_{\text{garb}}^{\prime}(1)$
[…]
$\lambda_{0}$
G.4 Principal’s Garbling Preferences
Similarly to Section D.4, the principal always prefers for more information to be revealed. In fact, the garbling parameter $\varepsilon$ directly interpolates between the concealed and revealed utilities for the principal.
@@ G.4 Principal’s Garbling Preferences
$\Pi_{\text{con}}(p^{*})\leq\Pi_{\text{garb}}(\varepsilon)\leq\Pi_{\text{rev}}(\rho^{*})$ for all $\varepsilon$.
Lemma 13 (Strict principal improvement).
Suppose $\theta=\frac{1}{2}$. Suppose the principal’s utility is strictly concave (Assumption 4). If the MLRP holds (Assumption 6), then $\Pi_{\text{garb}}^{\prime}(\varepsilon)>0$ for all $\varepsilon\in[0,1]$.
G.5 Garbling vs. Restricted Price Discrimination
Garbling is not the only way to interpolate between the full concealment and full revelation settings. In the price discrimination literature, there is an established model that interpolates between full price discrimination and no price discrimination by restricting that price difference between market segments can be no greater than some parameter $r$. Wright [1993] models this restriction as arising from a “cost of transport” or arbitrage between two markets. Aguirre et al. [2010] apply this interpolation by analyzing the marginal effect of $r$ on total welfare. We refer to this interpolation using $r$ as a restricted price discrimination model. We also leveraged this technique to analyze the agent’s utility in Section D.2.2.
Principal utility, $\Pi$ Refer to caption
Agent utility, $V$
Appendix H Proofs from Section G
Here we give proofs for results for the garbling model presented in Section G.
H.1 Proofs from Section G.1
@@ Appendix H Proofs from Section G
Suppose $\theta=\frac{1}{2}$. Suppose $F_{0},F_{1}$ are continuously differentiable CDFs, the principal’s utility is strictly concave (Assumption 4), and the MLRP holds (Assumption 6). Then $p_{0}^{\prime}(\varepsilon)>0$ and $p_{1}^{\prime}(\varepsilon)<0$ for all $\varepsilon\in[0,1]$.
@@
H.2 Proofs from Section 5.2
We first expand $V_{\text{garb}}(\varepsilon)$ for $\theta=\frac{1}{2}$.
@@
$+\frac{1}{2}\left((V_{0}(p_{0}(\varepsilon))-V_{0}(p_{1}(\varepsilon)))-(V_{1}(p_{0}(\varepsilon))-V_{1}(p_{1}(\varepsilon)))\right).$
@@
$+\frac{1}{2}\left(-\Delta(p_{0}(\varepsilon),p_{1}(\varepsilon))\right).$
@@
$-\Delta(p_{0}(\varepsilon),p_{1}(\varepsilon)).$
@@
$z_{x}(p)=\frac{V^{\prime}_{x}(p)}{\Pi_{x}^{{}^{\prime\prime}}(p)}.$
@@
H.3 Proofs from Section G.4
@@
$+\frac{1}{2}(b-p_{1}(\varepsilon))\left(\frac{1+\varepsilon}{2}F_{1}(p_{1}(\varepsilon))+\frac{1-\varepsilon}{2}F_{0}(p_{1}(\varepsilon))\right).$
@@
By the MLRP and strict concavity of $\Pi_{0}(p),\Pi_{1}(p)$, we have that $p_{1}(\varepsilon)<p_{0}(\varepsilon)<p_{0}^{*}$. Strict concavity of $\Pi_{0}(p)$ and the optimality of $p_{0}^{*}$ for $\Pi_{1}$ then implies that
@@
Appendix I Further Uber and Lyft Experiment Details
Here we provide additional details on the experiment setup and implementation using the Uber and Lyft dataset.
I.1 Scenario 2: Revelation to Refine an Existing Pricing Model.
As an additional scenario, we suppose that the principal starts with a slightly more sophisticated pricing model as a baseline. Suppose Uber decides to offer a non-negative price adjustment $p$ on top of their existing pricing model as an incentive for drivers to switch to their platform. The agent can choose to reveal $X=$ Lyft’s surge multiplier (which Uber is otherwise unable to observe), in which case Uber’s price adjustment would depend on $X$. To approximate Uber’s initial pricing model, we train a linear model targeting Uber’s price over all features in Uber’s dataset. [^5] A linear model is, of course, still far from Uber’s actual pricing methodology, but the purpose of this illustration is to have a model that depends on more features initially. Let $C$ be the difference between Uber’s estimated price (on the Lyft dataset) and Lyft’s price.
I.1.1 Results.
We present results with the same switching value $b$ as earlier. Figure 9 shows that the agent also prefers to reveal $X=$ Lyft’s surge multipler. Letting $Z^{t}$ be a binarized version of the surge multiplier, Figure 9 also shows that there exist at least one value of $t$ for which the agent prefers garbling over revelation. Since the surge multiplier is never less than 1 in the data, and the agent already prefers to reveal for $t=1$, we do not observe any thresholds for which the agent strictly prefers to conceal. However, the agent’s value for revelation still decays to 00 as the threshold increases to its maximum. In this case, the agent does not gain any value from additional garbling on top of revelation. Overall, this scenario showed that a cost-correlated feature like the surge multiplier was beneficial to an agent to reveal, even when the principal platform’s existing pricing model already depended on other features like distance.
Scenario 2: Agent’s utility difference for revealing surge multiplier
Figure 9: Differences between the agent’s revealed and concealed utilities in Scenario 2, with $X=$ Lyft’s surge multiplier. The labels are the same as Figure 4.
I.2 Estimating the Principal’s and Agent’s Utilities
For binarized features:
For a binarized feature $Z^{t}$, we directly estimate the conditional pdfs $f_{0}$ and $f_{1}$ using kernel density estimation over the data conditioned on $Z^{t}=0,1$ respectively. Specifically, we use the Scipy gaussian_kde method with default bandwidth Virtanen et al. [2020]. The principal and agents’ revealed values are built directly on $f_{0}$, $f_{1}$, and the empirical estimate for $\theta$. The hidden cost distribution is given by $f(c)=(1-\theta)f_{0}(c)+\theta*f_{1}(c)$.
For full features:
When estimating the agent’s value difference for revealing a full non-binarized feature $X$, we substitute the empirical expectation over the dataset for both the principal and the agent’s values. That is, for a dataset with points $\{(C_{i},X_{i})\}_{i=1}^{n}$, we estimate the principal’s revealed value as
$\Pi_{\text{rev}}(\rho)=\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho(X))(b-\rho(X))]\approx\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbbm{1}}(C_{i}<\rho(X_{i}))(b-\rho(X_{i})).$
Similarly, we estimate the agent’s revealed value as
$V_{\text{rev}}(\rho)=\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho(X))(\rho(X)-C)]\approx\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbbm{1}}(C_{i}<\rho(X_{i}))(\rho(X_{i})-C_{i}).$
The principal and agents’ hidden values are also estimated empirically using only the points $\{C_{i}\}_{i=1}^{n}$.
These estimates are then used to compute the optimal hidden price $p^{*}$, revealed price function $\rho^{*}$, and the agent’s value differences.
I.3 Parameterizing $\rho(x)$ for Continuous $X$
When the variable $X$ revealed is continuous, we parameterize the principal’s pricing function $\rho(x)$ as linear for simplicity and tractability. This is more restrictive than any non-parametric and arbitrarily expressive $\rho(x)$, though not uncommon in practice. In the experiments, the agent also operates under the assumption that $\rho(x)$ would be linear. For the purposes of our model, the most important assumption is that the agent is aware of the family of the pricing functions over which the principal is optimizing, $\rho\in\mathcal{F}$. The value difference for the agent between revealing and concealing will ultimately also depend on the family $\mathcal{F}$.
I.4 Garbling Amounts
For each $Z^{t}$, we compute the optimal garbling amount
$\varepsilon^{*}=\underset{\varepsilon\in[0,1]}{\arg\max}\;\;V_{\text{garb}}(\varepsilon).$
Notably, for the garbling distribution $\xi$, we set $\gamma=\theta=P(Z^{t}=1)$ for each $Z^{t}$. This ensures that the marginal distribution of the garbled variable $Y$ matches the marginal distribution of $Z^{t}$ for each $t$.
I.5 Scenario 2 Costs
To construct the cost variable $C$ for Scenario 2, we first estimate Uber’s pricing model by training an ordinary least squares linear regression model on Uber’s dataset using the price column as the target, and all features other than the surge multiplier as inputs. The $R^{2}$ score on the training data is $0.1131$. We use Scikit Learn’s linear_model function Pedregosa et al. [2011].
[…]
Note that the principal optimizing its price using this non-negative cost variable is equivalent to the principal solving the constrained optimization problem for price restricted to non-negative prices only.
I.6 Results for Different Values of $b$
Figures 4 and 9 show results when the principal’s value is $b=150$, or $1.5\bar{C}$, where $\bar{C}=100$. Figures 10 and 11 show the same analysis for $b\in\{50,100,200\}$.
Scenario 1: Agent’s utility difference for revealing distance
$b=200$
$b=100$
$b=50$
Figure 10: Differences between the agent’s revealed and concealed utilities in the Scenario from Section 6, with $X=$ distance, for different values of $b$. The labels are the same as Figure 4.
Scenario 2: Agent’s utility difference for revealing surge multiplier
$b=200$
$b=100$
$b=50$
Figure 11: Differences between the agent’s revealed and concealed utilities in Scenario 2, with $X=$ Lyft’s surge multiplier, for different values of $b$. The labels are the same as Figure 4.